### v2

# A Computational Social Simulation of Ageing and Care Accessibility in Italian Inner Areas

Abstract
Ageing societies face increasing strain on formal and informal care systems, particularly in low-density mountainous municipalities where sparse services and steep terrain constrain access. This study presents a spatially explicit agent-based model that integrates a road-network GIS, synthetic populations derived through Iterative Proportional Fitting, and behavioural heterogeneity to examine how alternative service configurations shape accessibility and caregiver burden. The model, applied to Premeno (Piedmont, Italy), compares a baseline distribution of ambulatory services with a relocation scenario at Villa Bernocchi. System-level indicators (Caregiver Effort, Overwhelm, Hours Not Cared, Walkability) and micro-spatial metrics (Walkability, Detour Ratio, Proximity) are analysed across 40 batches and 50 stochastic replications per scenario. Results reveal aggregate neutrality but pronounced local redistribution of accessibility. Sensitivity analysis shows that spatial impedance dominates accessibility, whereas behavioural capacity modulates care effort. The findings illustrate hallmark properties of complex adaptive social systems-emergence, heterogeneity, and feedback-demonstrating how computational social simulation can illuminate policy trade-offs between spatial efficiency, social equity, and care sustainability in ageing territories.
keywords:
accessibility, agent-based modelling, ageing, caregiver burden, computational social science, GIS
1 Introduction
Demographic ageing is transforming the structure and functioning of contemporary societies. As life expectancy rises and family networks shrink, the capacity of public and informal care systems to guarantee accessible and sustainable assistance becomes increasingly uncertain. Peripheral and mountainous areas are particularly exposed: dispersed settlement patterns, steep topography, and scarce public transport combine to amplify inequalities in access to health and social services. Understanding these dynamics requires analytical frameworks able to capture both the spatial heterogeneity of environments and the interactive behaviour of the individuals embedded within them. Computational social science provides such a framework by combining data-rich representations of social environments with agent-based simulation of micro-level decisions and interactions [1, 2]. Within this perspective, agent-based models (ABMs) have been applied to domains ranging from urban mobility and spatial systems [3, 6] to health behaviour [9] and ageing policy [11]. Recent work has further situated ABM within population health frameworks [10]. Yet few studies integrate geographic information systems (GIS), demographic microdata, and survey-derived behavioural attributes in a single, reproducible simulation environment. This article presents a spatially explicit agent-based model (ABM) that couples a road network GIS with a behaviourally heterogeneous population of elder-caregiver dyads. The model integrates municipal indicators and survey-based parameters via synthetic population techniques to evaluate how ambulatory service reconfiguration affects both system-level performance and micro-spatial accessibility patterns. The empirical case focuses on the municipality of Premeno (Piedmont, Italy) and compares a baseline configuration (Scenario 1) with a relocation to Villa Bernocchi (Scenario 2). The study aims to quantify aggregate differences, reveal sub-group and neighbourhood heterogeneity, and identify policy-relevant trade-offs.
Novel contribution. This study advances the field of computational social science by operationalising ageing-care accessibility as a complex adaptive social system. It unifies three methodological strands: (i) synthetic population generation through Iterative Proportional Fitting (IPF), (ii) GIS-embedded agent-based simulation of ageing and caregiving dynamics, and (iii) global sensitivity analysis for mechanism discovery. Together, these layers form a reproducible and data-rich framework that links the social-data synthesis tradition of population science with the mechanism-oriented exploration characteristic of computational social science. By integrating spatial, demographic, and behavioural dimensions, the model reveals how micro-level mobility and care-allocation behaviours give rise to macro-level accessibility and burden patterns, thereby supporting the evaluation of care sustainability in ageing territories.
2 Methods
2.1 Model overview
The model represents a population of dyads composed of older adults and their informal caregivers. Each dyad operates within a georeferenced environment derived from high-resolution spatial data, including topography, road networks, and service locations.
Table 1: ODD-D summary for the AWSim model, integrating details from the Supplementary Material on initialization, stochasticity, reproducibility, and submodel design.
ODD-D Element Summary of AWSim Implementation
Purpose To examine how spatial configuration of services, behavioural heterogeneity, and stochastic variability jointly influence accessibility, walkability, and caregiver burden in ageing populations of Italian inner-area municipalities. The model explores trade-offs between elder autonomy, caregiver sustainability, and service coverage efficiency.
Entities (1) Elder (patient) agents; (2) Informal caregiver agents; (3) Service facilities (ambulatory, pharmacy, etc.); (4) Road-network nodes representing the GIS environment.
State variables For elders: ageing stage, ADL/IADL ability, health severity, care hours required, and mobility radius. For caregivers: available time, income, support network, mobility mode, and effort index. For services: location, capacity, accessibility, and interaction score. Global variables include walkability (WKB), caregiver effort (CEI), hours not cared (HNC), and caregiver overwhelm (COI).
Temporal scale Discrete time; 1 tick = 1 hour, with a day subdivided into three sessions (morning, afternoon, evening). Simulations typically run for one month after a 30-iteration warm-up.
Spatial scale Municipal-level graph derived from GIS layers (OSM road network, Copernicus DEM). Nodes represent buildings or intersections, mean spacing $\approx$ 25 m. Each node stores elevation, slope, safety, fatigue, and pleasantness scores used in walkability computation.
Process overview (1) Update ageing stage and mortality; (2) Recompute caregiver availability and effort; (3) Select service session and compute metrics (e.g., travel time/distance, energy cost) (4) Update functional ability and recompute WKB, CEI, HNC, and COI indices; (5) Record KPIs and aggregate statistics. Time advances synchronously for all agents.
Design concepts Emergence: accessibility and burden patterns emerge from decentralized dyad decisions. Adaptation: agents adjust walking radius, assistance frequency, and mobility mode. Heterogeneity: agents differ by age, income, support, and mobility attributes. Stochasticity: Gaussian, uniform, and log-normal components govern care frequency, ageing, income, and session duration. Learning and objectives: bounded-rational decisions under limited information. Reproducibility: ensured through per-agent hashing-based random seeding.
Input data Demographic and socioeconomic indicators from ISTAT; GIS data (OSM roads, Copernicus DEM); Structured caregiver interviews; Empirical mobility datasets; Literature-based behavioural parameters. See ESM 1 (§5)
Initialization Synthetic population generated via iterative proportional fitting (IPF) and imputation using municipal marginals. Agents are assigned to residential nodes proportional to population density. Services are located according to scenario configuration. Warm-up length determined by Schruben-Singh-Tierney test ($\approx$ 30 ticks). Experiments executed with 40 independent replications and batch-means estimation.
Submodels (1) Walkability Index (WKB) - slope-, safety-, and pleasantness-weighted accessibility. (2) Caregiver Effort Index (CEI) - integrates income, support, and mobility burden. (3) Hours Not Cared (HNC) - unmet assistance due to time or income constraints. (4) Caregiver Overwhelm (COI) - identifies dyads exceeding effort thresholds. All indices are bounded, continuous, and computed from session-level outcomes.
A simulation run advances in discrete time steps representing days; and, at each step, dyads update their decisions according to local accessibility conditions, accumulated effort, and contextual constraints such as travel distance, slope, and terrain resistance. The model’s primary purpose is to explore how the spatial configuration of essential services and the characteristics of caregivers and elders jointly affect accessibility, caregiving burden, and the overall sustainability of local care systems.
The model follows the updated ODD-D protocol [12], ensuring full transparency and reproducibility. A condensed checklist of its core components is reported in Table 1, while the complete protocol-including purpose, entities, state variables, process scheduling, and design concepts-is provided in the Electronic Supplementary Material (ESM 1). All stochastic processes (agent initialization, tie assignment, random choice) are controlled by a fixed random seed, ensuring reproducibility across runs. Initialization, warm-up, and randomisation procedures are detailed in ESM 1 (§5-6), including the Schruben-Singh-Tierney test for warm-up length and the hashing-based RNG seeding scheme. The model was developed in NetLogo 6.4 with Python extension for batch execution and sensitivity analysis.
The combination of stage-level and dyad-level attributes enables heterogeneity in both physiological decline and caregiving capacity. Each elder-caregiver pair evolves according to interacting state variables: the elder’s ageing stage and mobility define care demand, while the caregiver’s availability, employment, and support network determine effective supply. Service locations act as environmental constraints linking behavioural rules to spatial structure, allowing system-level accessibility patterns to emerge endogenously from micro-level interactions. These parameterisations collectively ensure that simulated outcomes are consistent with the diversity observed in real-world ageing and caregiving trajectories.
2.2 Scenarios and experimental design
Premeno, located in the Verbano-Cusio-Ossola province of Piedmont (Italy), represents a typical mountainous inner-area municipality characterised by low population density and high ageing ratios (Table 2).
Table 2: Demographic context of the case-study municipality compared with the regional and national levels. Premeno exhibits a markedly aged population and high settlement compactness, representative of small mountainous inner-area communities.
Metric Premeno Piedmont Italy
Population 746 4.3 M 59.4 M
Urbanized (%) 96.4 91.1 91.0
Elderly (>>65 yr) (%) 39.0 37.1 32.0
Figure 1: Study area: municipality of Premeno (Piedmont, Italy), an inner area characterised by low service density and steep morphology. The inset map (top right) shows its geographical position in northern Italy. Source: OpenStreetMap [osm].
2.3 Synthetic population and data integration
A synthetic population was generated using small-area municipal indicators aligned with national aggregates and survey-derived joint constraints. Iterative Proportional Fitting (IPF) adjusted seed microdata from the ISTAT AVQ 2019 survey until the weighted marginals matched the 8MilaCensus 2011 distributions [23, 21]. Each synthetic record combines demographic, socioeconomic, and health attributes; dyads are then formed according to conditional caregiving probabilities from the AVQ survey. Residential assignment follows population-density weights on the road-network nodes. This procedure yields an empirically grounded yet anonymised micro-dataset reproducing the spatial and social structure of the case-study municipality.
All data used to construct the synthetic population are publicly available from ISTAT, OpenStreetMap [22], and the Copernicus Land Monitoring Service [20]. No individual-level personal data are processed. The synthetic dataset can be regenerated using the procedures described in the supplementary material.
Two configurations are analysed for Premeno: Scenario 1 (baseline distribution of ambulatory services) and Scenario 2 (relocation to the Villa Bernocchi site). Each scenario was parameterised using empirical and literature-based constants summarised in Table 4 in the ESM (§5.1). Each scenario is executed in 40 experimental batches with 50 independent replications per batch (2,000 runs per scenario). Key Performance Indicators (KPIs) include: Caregiver Effort, Caregiver Overwhelm, Hours Not Cared, and a Walkability Index. Micro-level indices include Walkability (WKB), Effective Detour Ratio (EDR), and Household Proximity Index (HPI). Statistical comparisons employ Welch $t$-tests with multiplicity adjustment where applicable and cluster-robust checks as reported in the scenario analysis.
3 Sensitivity Analysis and Robustness
A two-stage global sensitivity analysis was conducted following Morris [7] and Saltelli and Annoni [8]. Stage 1 applied the Morris elementary-effects method (10 trajectories, $\Delta=0.5$) for qualitative screening; Stage 2 employed Sobol variance decomposition with Saltelli sampling ($n=512$ samples per parameter, N=2  [^0] N=2\,000 total model evaluations). Sensitivity analyses were implemented through an automated hybrid pipeline, wherein NetLogo-via its Python extension-triggered a command-line shell to launch a Jupyter notebook executing the SALib routines. This modular configuration enabled seamless coupling between model execution and parameter sampling while enforcing identical random seeds across replications to guarantee methodological reproducibility.
Spatial impedance parameters (e.g., distance-decay and acceptable walking-time thresholds) strongly affected accessibility measures, while parameters related to behavioural capacity and mobility shaped care effort and unmet hours. Replication-based diagnostics showed narrowing confidence intervals with increasing replications and stable rank ordering of parameter influence, supporting numerical robustness (full details in the AWSim technical report referenced below).
4 Results and Discussion
4.1 Emergent spatial patterns of accessibility
Under Scenario 1, walkability is highly heterogeneous within the municipal boundary. Values in the dedicated walkability analysis range approximately between 12 and 47 with a mean near 27, consistent with fragmented connectivity on steep terrain and distance to services. Under Scenario 2, the mean Walkability Index increases to about 0.34 and the variance decreases, indicating improved overall accessibility and greater spatial uniformity. However, micro-level indicators reveal a trade-off: route indirectness (EDR) increases and proximity (HPI) decreases for some central households. The relocation thus yields a more walkable environment with slightly less direct access for particular neighbourhoods (see Table 4).
4.2 Behavioural and social heterogeneity
Behavioural heterogeneity mediates how spatial change translates into caregiving burden. Higher-mobility dyads maintain acceptable accessibility even in peripheral areas, while mobility-limited dyads experience accelerated burden once care-capacity thresholds are reached. In the Premeno experiment, clusters of high-burden dyads coincide with lower walkability and higher dependency, evidencing differentiated outcomes across ageing stages. These patterns illustrate how individual-level variation and local morphology interact to produce uneven but interpretable social outcomes. Stage-wise results (Table 4) show that walkability (WKB) improves consistently across all stages, while the effective detour ratio (EDR) rises by roughly +0.9 to +1.1, signalling a spatial trade-off between usability and route efficiency. Proximity (HPI) exhibits small and often non-significant declines, reflecting marginal losses for central households.
4.3 System-level and micro-level effects
System-level KPI comparisons between Scenario 1 and Scenario 2 do not yield statistically significant differences after correction: caregiver effort changes are small, caregiver overwhelm declines marginally, walkability changes are negligible at the aggregate level, and hours not cared display a modest increase when using cluster-robust checks as reported in the scenario analysis (Table 3). By contrast, micro-spatial indicators reveal clear redistribution of accessibility: Scenario 2 enhances walkability and spatial uniformity while slightly increasing route detours (Table 4).
Table 3: System-level KPIs (global means across replications) by scenario. Aggregate differences are small and not significant after multiplicity adjustment.
KPI Scenario 1 Scenario 2 $\boldsymbol{\Delta}$ (S2-S1) Significance
Caregiver Effort 0.16 0.17 +0.01 n.s. (Holm)
Hours Not Cared 1.86 2.16 +0.30 n.s. (Holm)
Caregiver Overwhelmed n.s. (see text)
Walkability (aggregate) $\approx 0$ n.s.; see micro-indices
Table 4: Stage-stratified effects on micro-indices by scenario. Scenario 2 increases the detour ratio (EDR) consistently; walkability (WKB) gains are uniform and robust; proximity (HPI) declines in some stages. Means are per stage; $p$-values are from stage-wise tests reported in the scenario analysis.
Stage Index S1 mean S2 mean $\boldsymbol{\Delta}$ (S2-S1) $p$-value Direction
1 EDR 1.8303 2.7419 +0.9116 0.000279 S2 >> S1
HPI 4.1958 4.3000 +0.1042 0.753918 n.s.
WKB 34.3289 34.8505 +0.5216 $10^{-6}$ S2 >> S1
2 EDR 1.8284 2.9466 +1.1182 0.000012 S2 >> S1
HPI 4.8708 4.0875 0.7833-0.7833 0.006894 S2 << S1
WKB 32.3078 32.8436 +0.5358 $10^{-6}$ S2 >> S1
3 EDR 1.7310 2.8449 +1.1139 0.000023 S2 >> S1
HPI 4.3833 4.0875 0.2958-0.2958 0.430495 n.s.
WKB 31.3211 31.8530 +0.5319 $10^{-6}$ S2 >> S1
4 EDR 1.8688 2.8328 +0.9640 0.000099 S2 >> S1
HPI 4.6625 4.1000 0.5625-0.5625 0.059223 n.s. (trend $\downarrow$)
WKB 29.3189 29.8289 +0.5100 $10^{-6}$ S2 >> S1
Figure 2: Comparison of ambulatory service locations in the two simulated configurations. Top: Scenario 1 (current distribution of ambulatory services) shows fragmented accessibility with Walkability Index values ranging from 12 to 47 (mean $\approx$ 27). Bottom: Scenario 2 (relocation to Villa Bernocchi) yields a more homogeneous spatial distribution and higher average walkability (mean $\approx$ 34), indicating improved accessibility for peripheral households while maintaining service coverage for central areas. Grey-scale points represent residential nodes weighted by walkability index (0-50). Source: OpenStreetMap [22], Simulation Data.
This divergence between macro-neutrality and micro-variation illustrates that spatial interventions reallocate relative advantage rather than uniformly improving welfare. Such heterogeneous effects, typical of complex adaptive systems, highlight the need for multi-scale evaluation when interpreting policy outcomes.
4.4 Sensitivity-driven insights
The sensitivity analysis clarifies mechanisms behind the scenario results. Spatial impedance parameters directly govern accessibility outcomes, while behavioural capacity parameters condition the translation of spatial gains into reduced burden. Consequently, spatial redesign alone may not yield measurable welfare improvements without parallel organisational or social-support measures (e.g., additional service days, mobility assistance, tele-care).
4.5 Model robustness and theoretical implications
These results resonate with broader theories of social resilience and welfare ecology, which frame care systems as adaptive socio-technical networks balancing efficiency and equity under resource constraints [17, 18, 19]. In this perspective, the redistribution of accessibility observed in the simulation mirrors real-world adaptive cycles-where structural interventions shift local equilibria without necessarily improving aggregate welfare. By explicitly modelling feedbacks among space, behaviour, and resources, the simulation formalises these resilience mechanisms within a computational social-science architecture.
Each scenario comprised forty experimental batches with fifty stochastic replications (2,000 total runs), yielding stable outcome distributions with narrow confidence intervals. Rank-order consistency across replications supports numerical robustness, while cluster-robust checks reduce the risk of spurious significance. From a theoretical perspective, the experiment exhibits three hallmark properties of computational social systems: emergence (macro accessibility patterns arise from decentralised decisions on a geospatial network), heterogeneity (behavioural and demographic diversity produces unequal yet interpretable outcomes), and feedback (changes in accessibility redistribute caregiver workload, which in turn modifies spatial demand). These features align with canonical characteristics of complex adaptive systems [5, 1, 4] and illustrate a micro-macro bridge consistent with computational social science.
5 Conclusions and Policy Implications
Service relocation in Premeno produces aggregate neutrality with local redistribution: system-level indicators remain statistically unchanged, whereas micro-spatial metrics reveal differentiated gains and losses. Behavioural heterogeneity mediates these effects, indicating that infrastructure redesign must be complemented by organisational and social-support interventions to achieve equitable burden reduction. Beyond the case study, the framework is transferable to other European small-municipality contexts where ageing, morphology, and service distribution interact. Its modular architecture enables comparative analyses and scenario testing for evidence-based policy in health-access planning.
Acknowledgments
The authors thank the municipal stakeholders for facilitating access to spatial information and providing assistance during site visits and documentation studies.
Funding and Compliance Statement
This research was conducted within the AGE-IT project, Spoke 5 “Care Sustainability in an Ageing Society,” co-funded by the European Union - NextGenerationEU through the Italian National Recovery and Resilience Plan (PNRR), Mission 4, Component 2, Investment 1.3, under the Extended Partnership PE8 “Conseguenze e sfide dell’invecchiamento” (Project Age-It - A Novel Public-Private Alliance to Generate Socioeconomic, Biomedical and Technological Solutions for an Inclusive Italian Ageing Society - Ageing Well in an Ageing Society; Grant Agreement No. AGE-IT-PE00000015; CUP: H43C22000840006).
The simulator employed in this study is used exclusively for research and experimental purposes in collaboration with municipal authorities. Data use complies with the General Data Protection Regulation (Reg. (EU) 2016/679) and with the licensing terms of the Italian National Institute of Statistics (ISTAT). The system falls under the research exemption of the EU Artificial Intelligence Act (Reg. (EU) 2024/1689) and is not classified as medical device software under Reg. (EU) 2017/745.
Disclaimer. The views and opinions expressed in this article are those of the author alone and do not necessarily reflect those of the AGE-IT consortium, its partners, or the funding authority.
Declarations
Ethics approval: Not applicable.
Consent to participate / publish: Not applicable.
Data availability: Simulation data and code are available from the corresponding author upon reasonable request. Public datasets (municipal indicators and spatial layers) are referenced in the technical documentation.
Competing interests: The authors declare no competing interests.
Author contributions: R. Garrone: model design, implementation, experiments, analysis, writing. Please note that Professor Bandini provided supervision at an earlier stage of the research project but was not involved in the conception, writing, or approval of the present manuscript. Her name was initially included in error and has been removed accordingly.
ORCID: Garrone: 0009-0005-7060-6774.
Supplementary Material
The Supplementary Material provides a complete technical companion to the AWSim model, offering extended methodological and computational details that complement the main manuscript. It follows the updated ODD-D protocol [12], ensuring transparency and reproducibility in agent-based model reporting. Specifically, the Supplementary Material includes:
  1. (a)
    Full ODD-D description: A comprehensive account of the model’s purpose, entities, state variables, spatial and temporal structures, process scheduling, and design concepts. The description follows the ODD-D formalism and integrates tables summarizing key attributes of elder, caregiver, and service agents, as well as schematic representations of submodels and data flows.
  2. (b)
    Stochasticity and randomization methods: Detailed discussion of random components affecting dyad selection, aging transitions, walking preferences, assistance hours, and income variability. The document specifies the use of NetLogo’s Mersenne Twister RNG, per-agent substreams, and a hashing-based seeding scheme for reproducibility.
  3. (c)
    Initialization, warm-up, and run design: A dedicated section outlines model initialization from synthetic populations generated via Iterative Proportional Fitting (IPF), warm-up length estimation using the Schruben-Singh-Tierney test with batch means, and replication design for statistical convergence.
  4. (d)
    Parameterisation tables: Complete listings of baseline and empirical parameters (e.g., mobility speeds, income levels, effort thresholds, and ageing transition rates), each with corresponding data sources (ISTAT, survey data, literature benchmarks). A summary of initialization parameters is provided in Table 5.
  5. (e)
    Sensitivity and uncertainty analysis: Extended results from global sensitivity analyses using Morris and Sobol-Saltelli methods, illustrating first-order and total-order effects of financial and mobility parameters on key indicators such as Walkability (WKB) and Caregiver Effort (CEI).
  6. (f)
    Scenario results and maps: Comparative figures showing baseline (Premeno) and relocation (Villa Bernocchi) scenarios, including heatmaps of walkability, service reachability, and caregiver burden across spatial and temporal scales.
All supplementary files-including the full ODD-D protocol, parameter tables, initialization datasets, and simulation code-are available from the corresponding author upon reasonable request. The documentation ensures that every modeling decision, random process, and calibration parameter can be independently verified and replicated.