### v4

# A Computational Social Simulation of Ageing and Care Accessibility in Italian Inner Areas

Abstract
Ageing societies face increasing strain on formal and informal care systems, particularly in low-density mountainous municipalities where sparse services and steep terrain constrain access. This study presents a spatially explicit agent-based model that integrates a road-network GIS, synthetic populations derived through Iterative Proportional Fitting, and behavioural heterogeneity to examine how alternative service configurations shape accessibility and caregiver burden. The model, applied to Premeno (Piedmont, Italy), compares a baseline distribution of ambulatory services with a relocation scenario at Villa Bernocchi. System-level indicators (Caregiver Effort, Overwhelmed Caregivers, Hours Not Cared, Walkability) and micro-spatial metrics (Walkability, Detour Ratio, Proximity) are analysed across 40 batches and 50 stochastic replications per scenario. Results reveal aggregate neutrality but pronounced local redistribution of accessibility. Sensitivity analysis shows that spatial impedance dominates accessibility, whereas behavioural capacity modulates care effort. The findings illustrate distinctive properties of complex adaptive social systems - emergence, heterogeneity, and feedback - demonstrating how computational social simulation can highlight policy trade-offs between spatial efficiency, social equity, and care sustainability in ageing territories.
Keywords: accessibility, agent-based modelling, ageing, caregiver burden, simulation modelling, GIS.
1 Introduction
Demographic ageing is transforming the structure and functioning of contemporary societies. As life expectancy rises and family networks shrink, the capacity of public and informal care systems to guarantee accessible and sustainable assistance becomes increasingly uncertain. Peripheral and mountainous areas are particularly exposed: dispersed settlement patterns, steep topography, and scarce public transport combine to amplify inequalities in access to health and social services. Understanding these dynamics requires analytical frameworks able to capture both the spatial heterogeneity of environments and the interactive behaviour of the individuals embedded within them. Computational social science provides such a framework by combining data-rich representations of social environments with agent-based simulation of micro-level decisions and interactions [4, 5]. Within this perspective, agent-based models (ABMs) have been applied to domains ranging from urban mobility and spatial systems [6, 13] to health behaviour [26] and ageing policy [2]. Recent work has further situated ABM within population health frameworks [24]. Yet few studies integrate geographic information systems (GIS), demographic microdata, and survey-derived behavioural attributes in a single, reproducible simulation environment. This article presents a spatially explicit agent-based model (ABM) that couples a road network GIS with a behaviourally heterogeneous population of elder-caregiver dyads. The model integrates municipal indicators and survey-based parameters via synthetic population techniques to evaluate how ambulatory service reconfiguration affects both system-level performance and micro-spatial accessibility patterns. The empirical case focuses on the municipality of Premeno (Piedmont, Italy) and compares a baseline configuration (Scenario 1) with a relocation of ambulatory services to Villa Bernocchi (Scenario 2), reflecting a policy option motivated by service rationalisation and cost-containment considerations. The study aims to quantify aggregate differences, reveal sub-group and neighbourhood heterogeneity, and identify policy-relevant trade-offs.
1.1 Positioning and Contribution
This study advances simulation-based analysis by operationalising ageing-care accessibility as a complex adaptive social system. Despite extensive research on ageing, accessibility, and service provision, a persistent gap remains in how care accessibility is conceptualised and analysed in sparsely populated and morphologically constrained territories. Much of the foundational and applied literature continues to operationalise accessibility through static indices, optimisation-based location models, or aggregate planning heuristics that prioritise average distance minimisation or coverage efficiency at the population level [12, 3, 6, 13]. While valuable for descriptive assessment and infrastructure benchmarking, these approaches typically abstract from individual-level behavioural constraints and implicitly assume representative users, fixed behavioural responses, and monotonic relationships between spatial efficiency and welfare outcomes [9, 20, 18, 22].
However, ageing-care systems exhibit the defining characteristics of complex socio-spatial systems, in which mobility limitations, caregiving capacity, and spatial structure interact non-linearly over time. Empirical and theoretical work in welfare geography, health services research, and computational social science shows that interventions improving aggregate accessibility indicators may simultaneously generate uneven redistributions of burden and opportunity across social groups and neighbourhoods [25, 1, 5, 4]. Such micro–macro decoupling is difficult to detect using aggregate or optimisation-based frameworks and remains underexplored in spatially explicit, agent-based analyses of care accessibility.
To address this gap, the present study unifies three methodological strands: (i) synthetic population generation through Iterative Proportional Fitting (IPF), (ii) GIS-embedded agent-based simulation of ageing and caregiving dynamics, and (iii) global sensitivity analysis for mechanism discovery. Together, these components form a reproducible and data-rich framework that links the social-data synthesis tradition of population science with the mechanism-oriented exploration characteristic of computational social science. By conceptualising accessibility and caregiver burden as emergent outcomes of decentralised, behaviourally heterogeneous interactions embedded in a real spatial network [14, 10], the model reveals a mechanism that is largely invisible to aggregate or optimisation-based approaches: cost-driven spatial reconfiguration can yield aggregate neutrality while redistributing accessibility and caregiver burden across neighbourhoods and social groups. The contribution of this work is therefore not a specific policy prescription, but a demonstration of how micro-level behavioural constraints and spatial structure decouple efficiency and equity at the system level, illustrating the analytical value of agent-based simulation for the study of ageing-care systems [26, 24, 11].
2 Methods
The empirical case is used as a controlled simulation environment designed to explore system behaviour under alternative service configurations, rather than as a basis for direct policy prescription or empirical forecasting. For clarity, we distinguish between different levels of the simulation design: a scenario refers to a structural configuration of the system (e.g., service location layout), a batch denotes a specific parameter setting within a scenario, and a run corresponds to a single stochastic replicate of a given batch.
2.1 Model overview
The model represents a population of dyads composed of older adults and their informal caregivers. Each dyad operates within a georeferenced environment derived from high-resolution spatial data, including topography, road networks, and service locations.
Table 1: ODD-D summary for the AWSim model, integrating details from the Supplementary Material on initialization, stochasticity, reproducibility, and submodel design.
ODD-D Element Summary of AWSim Implementation
Purpose To examine how spatial configuration of services, behavioural heterogeneity, and stochastic variability jointly influence accessibility, walkability, and caregiver burden in ageing populations of Italian inner-area municipalities. The model explores trade-offs between elder autonomy, caregiver sustainability, and service coverage efficiency.
Entities (1) Elder (patient) agents; (2) Informal caregiver agents; (3) Service facilities (ambulatory, pharmacy, etc.); (4) Road-network nodes representing the GIS environment.
State variables For elders: ageing stage, ADL/IADL ability, health severity, care hours required, and mobility radius. For caregivers: available time, income, support network, mobility mode, and effort index. For services: location, capacity, accessibility, and interaction score. Global variables include walkability (WKB), caregiver effort (CEI), hours not cared (HNC), and caregiver overwhelm (COI).
Temporal scale Discrete time; 1 tick = 1 hour, with a day subdivided into three sessions (morning, afternoon, evening). Simulations typically run for one month after a 30-iteration warm-up.
Spatial scale Municipal-level graph derived from GIS layers (OSM road network, Copernicus DEM). Nodes represent buildings or intersections, mean spacing $\approx$ 25 m. Each node stores elevation, slope, safety, fatigue, and pleasantness scores used in walkability computation.
Process overview (1) Update ageing stage and mortality; (2) Recompute caregiver availability and effort; (3) Select service session and compute metrics (e.g., travel time/distance, energy cost) (4) Update functional ability and recompute WKB, CEI, HNC, and COI indices; (5) Record KPIs and aggregate statistics. Time advances synchronously for all agents.
Design concepts Emergence: accessibility and burden patterns emerge from decentralized dyad decisions. Adaptation: agents adjust walking radius, assistance frequency, and mobility mode. Heterogeneity: agents differ by age, income, support, and mobility attributes. Stochasticity: Gaussian, uniform, and log-normal components govern care frequency, ageing, income, and session duration. Learning and objectives: bounded-rational decisions under limited information. Reproducibility: ensured through per-agent hashing-based random seeding.
Input data Demographic and socioeconomic indicators from ISTAT; GIS data (OSM roads, Copernicus DEM); Structured caregiver interviews; Empirical mobility datasets; Literature-based behavioural parameters. See Supplementary Material (§5)
Initialization Synthetic population generated via iterative proportional fitting (IPF) and imputation using municipal marginals. Agents are assigned to residential nodes proportional to population density. Services are located according to scenario configuration. Warm-up length determined by Schruben-Singh-Tierney test ($\approx$ 30 ticks). Experiments executed with 40 independent replications and batch-means estimation.
Submodels (1) Walkability Index (WKB) - slope-, safety-, and pleasantness-weighted accessibility. (2) Caregiver Effort Index (CEI) - integrates income, support, and mobility burden. (3) Hours Not Cared (HNC) - unmet assistance due to time or income constraints. (4) Caregiver Overwhelm (COI) - identifies dyads exceeding effort thresholds. All indices are bounded, continuous, and computed from session-level outcomes.
A simulation run advances in discrete time steps representing days; and, at each step, dyads update their decisions according to local accessibility conditions, accumulated effort, and contextual constraints such as travel distance, slope, and terrain resistance. The model’s primary purpose is to explore how the spatial configuration of essential services and the characteristics of caregivers and elders jointly affect accessibility, caregiving burden, and the overall sustainability of local care systems.
The model follows the updated ODD-D protocol [11], ensuring full transparency and reproducibility. A condensed checklist of its core components is reported in Table 1, while the complete protocol-including purpose, entities, state variables, process scheduling, and design concepts-is provided in the Supplementary Material. All stochastic processes (agent initialization, tie assignment, random choice) are controlled by a fixed random seed, ensuring reproducibility across runs. Initialization, warm-up, and randomisation procedures are detailed in the Supplementary Material (§5-6), including the Schruben-Singh-Tierney test for warm-up length and the hashing-based RNG seeding scheme. The model was developed in NetLogo 6.4 with Python extension for batch execution and sensitivity analysis.
The combination of stage-level and dyad-level attributes enables heterogeneity in both physiological decline and caregiving capacity. Each elder-caregiver pair evolves according to interacting state variables: the elder’s ageing stage and mobility define care demand, while the caregiver’s availability, employment, and support network determine effective supply. Service locations act as environmental constraints linking behavioural rules to spatial structure, allowing system-level accessibility patterns to emerge endogenously from micro-level interactions. These parameterisations collectively ensure that simulated outcomes are consistent with the diversity observed in real-world ageing and caregiving trajectories.
2.2 Scenarios and experimental design
Premeno, located in the Verbano-Cusio-Ossola province of Piedmont (Italy), represents a typical mountainous inner-area municipality characterised by low population density and high ageing ratios (Table 2).
Table 2: Demographic context of the case-study municipality compared with the regional and national levels. Premeno exhibits a markedly aged population and high settlement compactness, representative of small mountainous inner-area communities.
Metric Premeno Piedmont Italy
Population 746 4.3 M 59.4 M
Urbanized (%) 96.4 91.1 91.0
Elderly (>>65 yr) (%) 39.0 37.1 32.0
Two configurations are analysed for Premeno: Scenario 1 (baseline distribution of ambulatory services) and Scenario 2 (relocation to the Villa Bernocchi site). Each scenario was parameterised using empirical and literature-based constants summarised in Table S4 in the Supplementary Material (§5.1). Each scenario is executed in 40 experimental batches with 50 independent replications per batch (2,000 runs per scenario). Key Performance Indicators (KPIs) include: Caregiver Effort Index (CEI), Caregiver Overwhelmed (CO), Hours Not Cared (HNC), and a Walkability Index (WKB); see Table 3. Micro-level indices include Walkability (WKB), Effective Detour Ratio (EDR), and Household Proximity Index (HPI). Statistical comparisons employ Welch $t$-tests with multiplicity adjustment where applicable and cluster-robust checks as reported in the scenario analysis. In this study, accessibility denotes potential spatial opportunity to reach services, operationalised via the Walkability Index ($\mathrm{WKB}$), whereas realised care outcomes are captured by caregiver- and dyad-level indicators, including the Caregiver Effort Index ($\mathrm{CEI}$), Hours Not Cared ($\mathrm{HNC}$), and the Caregiver Overwhelmed indicator ($\mathrm{CO}$).
Figure 1: Study area: municipality of Premeno (Piedmont, Italy), an inner area characterised by low service density and steep morphology. The inset map (top right) shows its geographical position in northern Italy. Source: OpenStreetMap [21].
2.3 Synthetic population and data integration
A synthetic population was generated using small-area municipal indicators aligned with national aggregates and survey-derived joint constraints. Iterative Proportional Fitting (IPF) adjusted seed microdata from the ISTAT AVQ 2019 survey until the weighted marginals matched the 8MilaCensus 2011 distributions [16, 15]. Each synthetic record combines demographic, socioeconomic, and health attributes; dyads are then formed according to conditional caregiving probabilities from the AVQ survey. Residential assignment follows population-density weights on the road-network nodes. This procedure yields an empirically grounded yet anonymised micro-dataset reproducing the spatial and social structure of the case-study municipality.
All data used to construct the synthetic population are publicly available from ISTAT, OpenStreetMap [21], and the Copernicus Land Monitoring Service [8]. No individual-level personal data are processed. The synthetic dataset can be regenerated using the procedures described in the supplementary material.
3 Sensitivity Analysis and Robustness
A two-stage global sensitivity analysis was conducted following Morris [19] and Saltelli & Annoni [23]. Stage 1 applied the Morris elementary-effects method (10 trajectories, $\Delta=0.5$) for qualitative screening; Stage 2 employed Sobol variance decomposition with Saltelli sampling ($n=512$ samples per parameter, N=2  [^0] N=2\,000 total model evaluations). Sensitivity analyses were implemented through an automated hybrid pipeline, wherein NetLogo-via its Python extension-triggered a command-line shell to launch a Jupyter notebook executing the SALib routines. This modular configuration enabled seamless coupling between model execution and parameter sampling while enforcing identical random seeds across replications to guarantee methodological reproducibility.
Spatial impedance parameters (e.g., distance-decay and acceptable walking-time thresholds) strongly affected accessibility measures, while parameters related to behavioural capacity and mobility shaped care effort and unmet hours. Replication-based diagnostics showed narrowing confidence intervals with increasing replications and stable rank ordering of parameter influence, supporting numerical robustness (full details in the AWSim technical report referenced below).
Table 3: Key performance indicators (KPIs) and notation used throughout the study. Units refer to the model’s internal time and distance conventions (see Supplementary Material for implementation details).
Symbol Name Unit Interpretation Aggregation level
$\mathrm{WKB}$ Walkability Index score (0–50) Potential spatial accessibility to services via the road network under mobility constraints node / household / dyad; summarised to municipality
$\mathrm{CEI}$ Caregiver Effort Index effort score (0–1) Care effort required from the informal caregiver to meet elder needs given accessibility and constraints dyad; summarised by stage / municipality
$\mathrm{CO}$ Caregiver Overwhelmed number Count of the caregivers unable to provide more care; proxy for overload risk dyad; summarised by stage / municipality
$\mathrm{HNC}$ Hours Not Cared hours per period Unmet care demand (hours of required care not delivered) under current constraints dyad; summarised by stage / municipality
  • Notes. All KPIs are internally normalized before aggregation. $\mathrm{WKB}$, $\mathrm{CEI}$, and $\mathrm{CO}$ are dimensionless indices derived from time- and network-based quantities via monotone transformations, while $\mathrm{HNC}$ is expressed in physical time units but may be rescaled for comparability. Aggregated values are computed as population- or dyad-weighted means unless stated otherwise. Formal definitions and implementation details for $\mathrm{WKB}$, $\mathrm{CEI}$, $\mathrm{CO}$, and $\mathrm{HNC}$ are provided in the Supplementary Material (SM, Sections S2–S4).
Table 4: Micro-level spatial and network indices used for mechanism analysis. Indices are computed locally using network-based distances and routing.
Symbol Name Unit Interpretation Aggregation level
$\mathrm{WKB}$ Walkability Index score (0–50) Local potential to reach essential services on foot via the road network under mobility constraints node / household / dyad
$\mathrm{EDR}$ Effective Detour Ratio ratio ($\geq 1$) Inefficiency of network-based walking routes relative to straight-line distance node / household / dyad
$\mathrm{HPI}$ Household Proximity Index score (dimensionless) Network-based proximity of a household to nearby households or support nodes household / dyad
  • Notes. All indices use network-based distances; Euclidean distance is used only as a reference for $\mathrm{EDR}$. $\mathrm{WKB}$ values are internally normalised and discretised for reporting. Formal definitions and computational details are provided in the Supplementary Material (SM, Sections S2–S3).
3.1 Alternative Methodological Approaches
Several alternative methodological approaches could, in principle, be used to study ageing-care accessibility in spatially constrained territories. These approaches are briefly discussed here to clarify the rationale for adopting an agent-based modelling framework.
First, static accessibility metrics and GIS-based indicators—such as cumulative opportunity measures, gravity-based indices, or isochrone analyses—are widely used to evaluate spatial access to services [12, 9, 22]. While effective for benchmarking infrastructure coverage and identifying underserved areas, these methods treat accessibility as a fixed property of space and population averages. They do not capture how individual mobility limitations, caregiving capacity, or scheduling constraints interact dynamically over time, nor how accessibility gains translate into realised care under heterogeneous behavioural conditions [18, 20].
Second, optimisation-based location and allocation models—such as maximal coverage, $p$-median, or network flow formulations—can be employed to design service configurations that minimise average distance or maximise population coverage under resource constraints [3]. Although valuable for planning and normative assessment, such models typically assume representative users and deterministic demand, and they optimise system-level objectives by construction. Consequently, they are ill-suited to analysing behavioural adaptation, distributional effects, or the emergence of unintended inequalities following spatial interventions, particularly in ageing-care systems characterised by heterogeneous constraints [6, 13].
Third, regression-based and quasi-experimental approaches using observational data have been applied to assess associations between accessibility measures and health or care outcomes [1, 25]. However, these methods require strong assumptions about exogeneity and stable unit treatment values, and they are limited in their ability to explore counterfactual spatial configurations. This limitation is especially pronounced in small or peripheral municipalities, where data sparsity and the absence of natural experiments constrain causal inference.
In contrast, the agent-based approach adopted in this study is specifically designed to address these limitations. By representing accessibility and caregiver burden as emergent outcomes of decentralised interactions between heterogeneous agents embedded in a real spatial network, the model enables the exploration of non-linear dynamics, micro–macro decoupling, and distributional effects that are not accessible through aggregate, optimisation-based, or purely statistical frameworks [14, 5, 4]. The aim is therefore not to replace existing methods, but to complement them by providing a mechanism-oriented perspective on ageing-care accessibility in complex territorial settings, consistent with established principles of agent-based modelling and computational social science [10, 11, 26, 24].
4 Results and Discussion
This study contributes to computational social science by demonstrating how ageing-care accessibility emerges from the interaction between spatial morphology, behavioural heterogeneity, and decentralised decision-making. Rather than producing uniform improvements, spatial reconfiguration of services generates compensatory dynamics: aggregate indicators remain largely unchanged, while accessibility and caregiver burden are redistributed across neighbourhoods and social groups.
From a theoretical perspective, this result illustrates a key property of complex adaptive social systems: efficiency-oriented interventions can be system-neutral yet socially redistributive. The model shows that spatial efficiency and social equity are not monotonically aligned, as behavioural constraints condition how spatial opportunities translate into realised accessibility. This mechanism is largely invisible to representative- agent models or static accessibility metrics, but becomes evident when micro-level heterogeneity and spatial networks are modelled explicitly.
The study also highlights the importance of multi-scale evaluation. While system-level indicators suggest stability, micro-spatial and stage-stratified analyses reveal differentiated effects, underscoring the risk of relying exclusively on aggregate metrics in policy assessment. Agent-based simulation thus serves not as a forecasting tool, but as an explanatory device for identifying structural trade-offs and unintended consequences of policy interventions.
4.1 Emergent spatial patterns of accessibility
A central finding of the simulation is the coexistence of aggregate neutrality with pronounced local and group-level redistribution of accessibility and caregiving burden. While system-level indicators show limited change across scenarios, micro-spatial and stage-stratified metrics reveal differentiated gains and losses, indicating a divergence between macro-level stability and micro-level reallocation. Under Scenario 1, walkability is highly heterogeneous within the municipal boundary. Values in the dedicated walkability analysis range approximately between 12 and 47 with a mean near 27, consistent with fragmented connectivity on steep terrain and distance to services. Under Scenario 2, the mean Walkability Index increases to about 34 and the variance decreases, indicating improved overall accessibility and greater spatial uniformity. However, micro-level indicators reveal a trade-off: route indirectness (EDR) increases and proximity (HPI) decreases for some central households. The relocation thus yields a more walkable environment with slightly less direct access for particular neighbourhoods (see Table 5). Such micro–macro divergence is a characteristic feature of agent-based representations of social systems, where aggregate stability can coexist with substantial local redistribution driven by heterogeneous constraints and interactions [5, 4].
Table 5: Stage-stratified effects on micro-indices by scenario. Scenario 2 increases the detour ratio (EDR) consistently; walkability (WKB) gains are uniform and robust; proximity (HPI) declines in some stages. Means are per stage; $p$-values are from stage-wise tests reported in the scenario analysis.
Stage Index S1 mean S2 mean $\boldsymbol{\Delta}$ (S2-S1) $p$-value Direction
1 EDR 1.8303 2.7419 +0.9116 0.000279 S2 >> S1
HPI 4.1958 4.3000 +0.1042 0.753918 n.s.
WKB 34.3289 34.8505 +0.5216 $10^{-6}$ S2 >> S1
2 EDR 1.8284 2.9466 +1.1182 0.000012 S2 >> S1
HPI 4.8708 4.0875 0.7833-0.7833 0.006894 S2 << S1
WKB 32.3078 32.8436 +0.5358 $10^{-6}$ S2 >> S1
3 EDR 1.7310 2.8449 +1.1139 0.000023 S2 >> S1
HPI 4.3833 4.0875 0.2958-0.2958 0.430495 n.s.
WKB 31.3211 31.8530 +0.5319 $10^{-6}$ S2 >> S1
4 EDR 1.8688 2.8328 +0.9640 0.000099 S2 >> S1
HPI 4.6625 4.1000 0.5625-0.5625 0.059223 n.s. (trend $\downarrow$)
WKB 29.3189 29.8289 +0.5100 $10^{-6}$ S2 >> S1
4.2 Behavioural and social heterogeneity
The redistribution observed across scenarios is not solely a spatial effect but is strongly conditioned by behavioural heterogeneity among elder–caregiver dyads. Differences in mobility, caregiving capacity, income, and support networks shape how spatial accessibility translates into realised care outcomes. Behavioural heterogeneity mediates how spatial change translates into caregiving burden. Higher-mobility dyads maintain acceptable accessibility even in peripheral areas, while mobility-limited dyads experience accelerated burden once care-capacity thresholds are reached. In the Premeno experiment, clusters of high-burden dyads coincide with lower walkability and higher dependency, evidencing differentiated outcomes across ageing stages. These patterns illustrate how individual-level variation and local morphology interact to produce uneven but interpretable social outcomes. Stage-wise results (Table 5) show that walkability (WKB) improves consistently across all stages, while the effective detour ratio (EDR) rises by roughly +0.9 to +1.1, signalling a spatial trade-off between usability and route efficiency. Proximity (HPI) exhibits small and often non-significant declines, reflecting marginal losses for central households. These findings reinforce the role of behavioural heterogeneity as a conditioning mechanism that filters spatial opportunity, a dynamic that cannot be represented by representative-agent or purely spatial accessibility models [10, 26].
4.3 System-level and micro-level effects
The relocation scenario considered in this study reflects a cost-containment and service-rationalisation logic frequently discussed in small-municipality health planning. Rather than assuming accessibility improvement as an objective, the scenario provides a test case for examining the distributive consequences of efficiency-driven spatial reconfiguration. System-level KPI comparisons between Scenario 1 and Scenario 2 do not yield statistically significant differences after correction: caregiver effort changes are small, caregiver overwhelm declines marginally, walkability changes are negligible at the aggregate level, and hours not cared display a modest increase when using cluster-robust checks as reported in the scenario analysis (Table 6). By contrast, micro-spatial indicators reveal clear redistribution of accessibility: Scenario 2 enhances walkability and spatial uniformity while slightly increasing route detours (Table 5). From this perspective, the simulation highlights how efficiency-oriented interventions may reallocate disadvantage without producing aggregate welfare gains, a pattern consistent with broader discussions of resilience and unintended effects in public health and welfare systems [7, 27].
Table 6: System-level KPIs (global means across replications) by scenario. Aggregate differences are small and not significant after multiplicity adjustment.
KPI Scenario 1 Scenario 2 $\boldsymbol{\Delta}$ (S2-S1) Significance
Caregiver Effort 0.16 0.17 +0.01 n.s. (Holm)
Hours Not Cared 1.86 2.16 +0.30 n.s. (Holm)
Caregiver Overwhelmed n.s. (see text)
Walkability (aggregate) $\approx 0$ n.s.; see micro-indices
As with all agent-based models, the present simulation is intended as an explanatory and exploratory device rather than a predictive instrument. Its scope is bounded by stylised behavioural rules, exogenously specified institutional configurations, and simplified representations of health and care trajectories. This divergence between macro-neutrality and micro-variation illustrates that spatial interventions reallocate relative advantage rather than uniformly improving welfare. Such heterogeneous effects, typical of complex adaptive systems, highlight the need for multi-scale evaluation when interpreting policy outcomes. Explicitly acknowledging these boundary conditions clarifies the class of questions the model is designed to address and aligns with established best practices for transparent and reproducible agent-based modelling [11, 5]. Accordingly, validation is approached in an exploratory and structural sense, focusing on the internal coherence of mechanisms, the plausibility of agent behaviours, and the robustness of observed patterns under parameter variation. Rather than seeking point-wise empirical prediction, the model is intended to assess whether hypothesised interactions between spatial structure, behavioural constraints, and service configuration are sufficient to generate the observed qualitative patterns of accessibility and caregiver burden. Global sensitivity analysis is therefore used as a primary validation instrument to identify influential parameters and to distinguish structural effects from artefacts of parameterisation. This approach is consistent with established practices in agent-based and simulation modelling when the aim is to explore mechanisms and trade-offs in complex social systems.
Figure 2: Comparison of ambulatory service locations in the two simulated configurations. Top: Scenario 1 (current distribution of ambulatory services) shows fragmented accessibility with Walkability Index values ranging from 12 to 47 (mean $\approx$ 27). Bottom: Scenario 2 (relocation to Villa Bernocchi) exhibits a more homogeneous spatial distribution and higher average walkability values (mean $\approx$ 34), relative to Scenario 1, while maintaining service coverage for central areas. Grey-scale points represent residential nodes weighted by Walkability Index (0–50). Values shown are node-level WKB scores, aggregated over runs as means. Source: OpenStreetMap [21], simulation data.
4.4 Sensitivity-driven insights
Beyond the specific case study, the results have implications for how ageing-care systems are conceptualised and analysed within simulation-based modelling and analysis. The model demonstrates how spatial morphology, behavioural constraints, and decentralised decision-making jointly shape emergent accessibility and burden patterns. The sensitivity analysis clarifies mechanisms behind the scenario results. Spatial impedance parameters directly govern accessibility outcomes, while behavioural capacity parameters condition the translation of spatial gains into reduced burden. Consequently, spatial redesign alone may not yield measurable welfare improvements without parallel organisational or social-support measures (e.g., additional service days, mobility assistance, tele-care). More generally, this study illustrates how agent-based simulation can be used to formalise structural trade-offs in ageing-care systems by linking micro-level behaviour, spatial structure, and emergent system-level outcomes, consistent with core principles of computational social science and complex adaptive systems theory [4, 14].
4.5 Emergent socioeconomic and caregiving regimes
While aggregate indicators show limited net change between scenarios, the simulation outputs exhibit pronounced heterogeneity at the dyad level. To characterise this heterogeneity, a post-simulation clustering analysis was conducted on aggregated outputs across all experimental batches. The objective is not prediction, but interpretation: to identify recurrent socioeconomic and behavioural regimes through which spatial accessibility and caregiving burden are differentially realised.
Clustering was applied to a standardised feature space combining structural socioeconomic attributes (income, education, employment, job skill, car ownership) and caregiving outcomes (hours of care, caregiver effort, unmet care, and caregiver overwhelm). The number of clusters was selected using concordant internal validity criteria (Silhouette, Davies–Bouldin, and Calinski–Harabasz indices), which consistently supported a four-cluster solution. Alternative clustering formulations (Gaussian Mixture Models and density-based methods) yielded structurally comparable partitions, indicating that the identified regimes are robust to algorithmic choice.
Table 7 summarises the four emergent regimes using cluster-level means of key indicators and qualitative semantic labels derived from their relative positioning with respect to global averages.
Table 7: Emergent socioeconomic and caregiving regimes identified through clustering of simulation outputs. Labels reflect relative positions with respect to global means of socioeconomic endowment and caregiving indicators.
Cluster Socioeconomic profile Care profile Interpretation
C0 Predominantly lower SES profiles, including a subset of elderly living alone without an active caregiver Low recorded caregiving effort, negligible care hours, no reported burden Low observed burden arising from a combination of self-managed ageing and limited caregiver activation
C1 High income, high education, high job skill, high mobility High effort, long care hours, highest caregiver overwhelm Resource-rich but overloaded: formalised or intensive care arrangements amplify burden despite favourable SES
C2 Low income, lower education, limited job skill High effort, long care hours, moderate burden Resilient but vulnerable: high time investment compensates for scarce resources, masking future burnout risk
C3 Mixed SES, gender-imbalanced composition High perceived effort, short care hours, low overwhelm Emotionally intensive care: low time commitment but high subjective strain
Several insights follow. First, caregiver burden is not monotonically associated with socioeconomic disadvantage. The cluster exhibiting the highest overload is socioeconomically advantaged (C1), suggesting that resource availability may enable more demanding or formalised care trajectories that intensify time pressure and psychological strain. Second, low observed burden does not necessarily imply favourable conditions: clusters C0 and C3 combine low overload with indicators of latent constraint or emotional stress, which would be invisible in aggregate statistics. Third, the coexistence of C1 and C2 highlights a key decoupling between time investment and perceived burden, reinforcing the interpretation of caregiving as a multidimensional process shaped by both material resources and behavioural adaptation. The interpretation of C0 requires particular care. This cluster includes a heterogeneous set of agents, among which elderly individuals living alone and self-managing daily needs are represented alongside dyads with minimal caregiving activation. Low observed burden in this regime therefore does not necessarily indicate favourable accessibility conditions, but rather reflects limited engagement with caregiving processes—either because needs remain below intervention thresholds or because care is managed autonomously. As a result, C0 remains weakly coupled to both spatial accessibility improvements and caregiving redistribution mechanisms. This distinguishes it structurally from caregiver-mediated regimes (C1–C3), which respond more directly to changes in service configuration.
These regimes provide a behavioural interpretation layer for the scenario results. Aggregate neutrality between service configurations conceals redistribution across clusters: spatial interventions reallocate accessibility and effort unevenly across socioeconomic profiles rather than improving outcomes uniformly. In this sense, clustering complements the micro-spatial analysis by identifying who gains or loses under structural change, not merely where changes occur. This reinforces the interpretation of the ageing-care system as a complex adaptive system in which efficiency-oriented spatial reconfiguration redistributes burden across heterogeneous agent groups without necessarily improving aggregate welfare.
4.6 Model robustness and theoretical implications
These results resonate with broader theories of social resilience and welfare ecology, which frame care systems as adaptive socio-technical networks balancing efficiency and equity under resource constraints [7, 27, 17]. In this perspective, the redistribution of accessibility observed in the simulation mirrors real-world adaptive cycles-where structural interventions shift local equilibria without necessarily improving aggregate welfare. By explicitly modelling feedbacks among space, behaviour, and resources, the simulation formalises these resilience mechanisms within a computational social-science architecture.
Each scenario comprised forty experimental batches with fifty stochastic replications (2,000 total runs), yielding stable outcome distributions with narrow confidence intervals. Rank-order consistency across replications supports numerical robustness, while cluster-robust checks reduce the risk of spurious significance. From a theoretical perspective, the experiment exhibits three hallmark properties of computational social systems: emergence (macro accessibility patterns arise from decentralised decisions on a geospatial network), heterogeneity (behavioural and demographic diversity produces unequal yet interpretable outcomes), and feedback (changes in accessibility redistribute caregiver workload, which in turn modifies spatial demand). These features align with canonical characteristics of complex adaptive systems [14, 4, 10] and illustrate a micro-macro bridge consistent with computational social science.
4.7 Limitations and Future Work
Several limitations delimit the scope of the present work. Behavioural rules are stylised, institutional dynamics are exogenously specified, and health trajectories are simplified. The model does not capture endogenous policy adaptation, learning by institutions, or long-term demographic feedbacks. Accordingly, results should be interpreted as structural insights rather than empirical predictions.
Future research may extend the framework in three directions. First, incorporating adaptive policy agents would allow exploration of endogenous governance dynamics. Second, coupling the model with longitudinal empirical data could support calibration of behavioural rules and validation of emergent patterns. Third, comparative applications across multiple municipalities would enable systematic investigation of how morphology, demography, and service density jointly shape accessibility regimes.
5 Conclusions and Policy Implications
By making explicit the mechanisms through which spatial and behavioural factors interact, the study demonstrates how agent-based modelling can inform ageing and care policy by revealing not only whether interventions work on average, but for whom and where their consequences are most pronounced. Service relocation in Premeno produces aggregate neutrality with local redistribution: system-level indicators remain statistically unchanged, whereas micro-spatial metrics reveal differentiated gains and losses. Behavioural heterogeneity mediates these effects, indicating that infrastructure redesign must be complemented by organisational and social-support interventions to achieve equitable burden reduction. Beyond the case study, the framework is transferable to other European small-municipality contexts where ageing, morphology, and service distribution interact. Its modular architecture enables comparative analyses and scenario testing for evidence-based policy in health-access planning.
Acknowledgments
The authors thank the municipal stakeholders for facilitating access to spatial information and providing assistance during site visits and documentation studies.
Funding and Compliance Statement
This research was conducted within the AGE-IT project, Spoke 5 “Care Sustainability in an Ageing Society,” co-funded by the European Union - NextGenerationEU through the Italian National Recovery and Resilience Plan (PNRR), Mission 4, Component 2, Investment 1.3, under the Extended Partnership PE8 “Conseguenze e sfide dell’invecchiamento” (Project Age-It - A Novel Public-Private Alliance to Generate Socioeconomic, Biomedical and Technological Solutions for an Inclusive Italian Ageing Society - Ageing Well in an Ageing Society; Grant Agreement No. AGE-IT-PE00000015; CUP: H43C22000840006).
The simulator employed in this study is used exclusively for research and experimental purposes in collaboration with municipal authorities. Data use complies with the General Data Protection Regulation (Reg. (EU) 2016/679) and with the licensing terms of the Italian National Institute of Statistics (ISTAT). The system falls under the research exemption of the EU Artificial Intelligence Act (Reg. (EU) 2024/1689) and is not classified as medical device software under Reg. (EU) 2017/745.
Disclaimer. The views and opinions expressed in this article are those of the author alone and do not necessarily reflect those of the AGE-IT consortium, its partners, or the funding authority.
Declarations
Ethics approval: Not applicable.
Consent to participate / publish: Not applicable.
Data availability: Simulation data and code are available from the corresponding author upon reasonable request. Public datasets (municipal indicators and spatial layers) are referenced in the technical documentation.
Competing interests: The authors declare no competing interests.
AI: During the preparation of this work, the author used ChatGPT (OpenAI) to assist with language editing, clarity improvement, and formatting of the manuscript. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
ORCID: Garrone: 0009-0005-7060-6774.
Supplementary Material
The Supplementary Material provides a complete technical companion to the AWSim model, offering extended methodological and computational details that complement the main manuscript. It follows the updated ODD-D protocol [11], ensuring transparency and reproducibility in agent-based model reporting. Specifically, the Supplementary Material includes:
  1. (a)
    Full ODD-D description: A comprehensive account of the model’s purpose, entities, state variables, spatial and temporal structures, process scheduling, and design concepts. The description follows the ODD-D formalism and integrates tables summarizing key attributes of elder, caregiver, and service agents, as well as schematic representations of submodels and data flows.
  2. (b)
    Stochasticity and randomization methods: Detailed discussion of random components affecting dyad selection, aging transitions, walking preferences, assistance hours, and income variability. The document specifies the use of NetLogo’s Mersenne Twister RNG, per-agent substreams, and a hashing-based seeding scheme for reproducibility.
  3. (c)
    Initialization, warm-up, and run design: A dedicated section outlines model initialization from synthetic populations generated via Iterative Proportional Fitting (IPF), warm-up length estimation using the Schruben-Singh-Tierney test with batch means, and replication design for statistical convergence.
  4. (d)
    Parameterisation tables: Complete listings of baseline and empirical parameters (e.g., mobility speeds, income levels, effort thresholds, and ageing transition rates), each with corresponding data sources (ISTAT, survey data, literature benchmarks). A summary of initialization parameters is provided in Table 5.
  5. (e)
    Sensitivity and uncertainty analysis: Extended results from global sensitivity analyses using Morris and Sobol-Saltelli methods, illustrating first-order and total-order effects of financial and mobility parameters on key indicators such as Walkability (WKB) and Caregiver Effort (CEI).
  6. (f)
    Scenario results and maps: Comparative figures showing baseline (Premeno) and relocation (Villa Bernocchi) scenarios, including heatmaps of walkability, service reachability, and caregiver burden across spatial and temporal scales.
All supplementary files-including the full ODD-D protocol, parameter tables, initialization datasets, and simulation code-are available from the corresponding author upon reasonable request. The documentation ensures that every modeling decision, random process, and calibration parameter can be independently verified and replicated.